If 535 L of gas at 25.0C is heated to 75 0C  at constant pressure, calculate the new volume of the sample.


1
Expert's answer
2021-01-20T03:17:35-0500

pV= nRT. n=pV/RT = 101.325 x 535/8.314 x 298 = 21,8798384 moles.

V= nRT/p = 21,8798384 x 8.314 x 348/101.325 = 624.765 L.